OpenStudy (anonymous):

help with square root problem?? sqrt 5p+5 = sqrt 6p+1

OpenStudy (anonymous):

\[\sqrt{5p+5}=\sqrt{6p+1}\]

OpenStudy (anonymous):

@amistre64 please help?

OpenStudy (amistre64):

the under parts have to be equal so the radical is superfluous

OpenStudy (amistre64):

5p+5 = 6p+1

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

what about \[\sqrt{2w+5}+4=\sqrt{25}\]

OpenStudy (amistre64):

well, do you know the sqrt of 25?

OpenStudy (anonymous):

yeah 5 its the 2w that confuses me

OpenStudy (amistre64):

forget teh 2w, lets work it for what it is: \[\sqrt{2w+5}+4=\sqrt{25}\] \[\sqrt{2w+5}+4=5\] \[\sqrt{2w+5}+4-4=5-4\] \[\sqrt{2w+5}=1\]

OpenStudy (amistre64):

we can either square each side or just recall that the sqrt(1) = 1

OpenStudy (amistre64):

2w+5 = 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so w=-2

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you! you dont have to, but would you want to help me with 3 more problems?

OpenStudy (amistre64):

if got like 20 other people to see to, post them in a new question and see who bites :)